the displacement of a particle varies with time as s=3t square+2t find the velocity of the particle at t=1sec

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Akash
 
s = 3 t² + 2t
Now
V = ds/dt = 6t +2
Hence velocity at t = 1 sec
V = 6 +2 = 8 m/sec
 
Regards
Arun (askIITians forum expert)